OpenStudy (anonymous):

Which graph represents the inequality five times the quantity seven plus twelve x end quantity minus thirteen is less than or equal to negative eight?

OpenStudy (anonymous):

@calculusxy

OpenStudy (abb0t):

E

OpenStudy (calculusxy):

Can you provide the graph please?

OpenStudy (anonymous):

a number line shows a closed dot at negative two and an arrow extending from that dot pointing toward the right a number line shows a closed dot at negative two and an arrow extending from that dot pointing toward the left a number line shows a closed dot at negative one-half and an arrow extending from that dot pointing toward the right number line shows a closed dot at negative one-half and an arrow extending from that dot pointing toward the left

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

\[5(7+12x) -13 \le -8\] \[35 +60x -13 \le -8\] \[60x \le -30\] \[x \le -\frac{1}{2}\]